OpenStudy (anonymous):

intergration problem, how to do intergration of [3sin^2(x/2)]^2

OpenStudy (amistre64):

that is a mess aint it

OpenStudy (amistre64):

9 sin^4(x/2) ; and pull out the 9 for later \(\int\) sin^4(x/2) dx

OpenStudy (anonymous):

eh got a typo should be ∫ (3sin(x/2))^2 dx

OpenStudy (anonymous):

sorry dude

OpenStudy (amistre64):

sin^2 sin^2 ; where sin^2 = 1 -cos^2 sin^2(1-cos^2) \(\int\) sin^2 - sin^2cos^2

OpenStudy (amistre64):

........ ok

OpenStudy (amistre64):

(3 sin(x/2))^2 = 9 sin^2(x/2) ; pull out the 9 9 \(\int\) sin^2(x/2) dx 9 \(\int\) \(\frac{2}{2}\)[sin(x/2)]^2 dx ; pull out the 2 on top 9.2 \(\int\frac{1}{2}\) [sin(x/2)]^2 dx

OpenStudy (amistre64):

im gonna have ta write it on paper to keep track of it; typing and mathing dont go together :)

OpenStudy (anonymous):

lol this is hard

OpenStudy (anonymous):

Ok, here is how to do it. First, do a substitution u=x/2 - just to make it a bit easier to write Then, realize that sin^2=(1-cos(2x))/2 I think you can get it from there...

OpenStudy (amistre64):

yeah, i was just recalling that ..... lol thnx

OpenStudy (anonymous):

i look at the answer and its says cos 2x = 1+2sin^2x gives sin^2x = (1-cos2x)/2 cos x = 1+ 2sin^2(x/2) gives sin^2(x/2) = (1-cos x)/2 MESSY

OpenStudy (amistre64):

\[sin^2(x/2)=\frac{1}{2}-\frac{cos(x)}{2}\]

OpenStudy (anonymous):

ummm i dont get the subsitution

OpenStudy (anonymous):

Oh, yeah, I suppose you don't need it.

OpenStudy (amistre64):

its really about the trib identity for cos(2x) = 1 - 2sin^2(x)

OpenStudy (amistre64):

err... trig even lol

OpenStudy (amistre64):

\[9\int\left(\frac{1}{2}-\frac{1}{2}cos(x)\right)\ dx\]

OpenStudy (anonymous):

ummm thx but can u explain in the u subsitution way? coz i always get stuck in this method

OpenStudy (amistre64):

there is no usub method for this .... its jusst noticinng that sin^2 is an identity and this is how you integrate it

OpenStudy (anonymous):

Well...I suppose you could attempt to come up with a creative u, likely involving inverse trig functions, which will be integrable by basic methods, but would probably be a pain to write...

OpenStudy (anonymous):

lol ok thx both of u ^^

OpenStudy (amistre64):

youre welcome :)